OpenStudy (anonymous):

Hey everyone, we're learning line intercept methods to graph lines , my question is: how can I find the x intercept in x - y = 4 Many thanks ^^

OpenStudy (anonymous):

\[ax+by+c=0\] be a line.... \[ax+by=-c\] \[\frac{ax}{-c}+\frac{by}{-c}=1\] \[\frac{x}{\frac{a}{-c}}+\frac{y}{\frac{a}{-c}}=1\] the the thing below X is x intercept..

OpenStudy (anonymous):

try for the given line....in your case you should start with the second equation...

OpenStudy (anonymous):

I don't understand what you wrote, can you give me a solution using my equation please

OpenStudy (callisto):

To find x-intercept, put y=0 into the equation x - y = 4. x - (0) = 4 Solve x.

OpenStudy (anonymous):

Oh ok now I understand how this works, thank you so much :)

OpenStudy (callisto):

Welcome!